"I'm thinking of a number that, when you multiply it by 9, it is the same as adding 8 to 7 times the number. What is the number?"

Found 2 solutions by Jk22, stanbon:
Answer by Jk22(389)

Let n be this number, then 9n = 8+7n

hence 2n=8, this number n is 4, indeed :

9*4=36 and 7*4=28 28 and 8 is 36.

Answer by stanbon(75887)

"I'm thinking of a number that, when you multiply it by 9, it is the same as adding 8 to 7 times the number. What is the number?"
------------------------------------------------
Let the number be "x":
Equation:
9x = 8 + 7x
-----
2x = 8
---
x = 4
=============
Cheers,
Stan H.